Table Manufacturing Company produces one style of tables.

the following data pertain to producing one table

Planned production/month

?

units (one table)

50

Piece of woods (M)

20

Estimated M price

$20

Actual production

?

Quantity purchased (QP) from M

19

Actual price (AP)

$21

Material?volume (efficiency)?variance?

?

$19, Favorable

$19, Unfavorable

$20, Favorable

$20, Unfavorable

Given: Total direct material variance $200 favorable (F), Direct material price variance $900 unfavorable (UF), Standard price $10 Standard quantities 750 Find actual price

1.

Less than $10

2.

Between $10 and $10.50

3.

Between $11.00 and $11.50

4.

More than $12.00
Answer #1

Given total material variance=200 (F)

SQ*SP- AQ*AP = $200

750*$10 - AQ*AP = $200

AQ*AP = $7300............(1)

Direct material price variance = $900 (UF)

AQ*(SP - AP). = $900(UF)

AQ*10 - AQ*AP = $900(UF)

AQ*10 - 7300 =$900(UF)

AQ =(7300-900)/10

  AQ = 640 tables

From (1). AQ*AP =$7300

AP =$7300/640

  AP = $ 11.41